OpenStudy (theviper):

A student walks from his house at 4 km/hr & reaches his school late by 5 minutes.If his speed has been 5 km/hr then he would have reached 10 minutes early.The distance of the school from his house??

OpenStudy (theviper):

@mathslover @Callisto @lgbasallote Plz help:)

mathslover (mathslover):

sorry i am busy but may be this will help u http://www.bankexam.co.in/uploads/exam_features/1008091281348289Time%20&%20Distance.pdf

OpenStudy (callisto):

Let d be the distance and t be the time spent if he can arrive school punctually. Formula used: speed = distance / time ''A student walks from his house at 4 km/hr & reaches his school late by 5 minutes'' 5 minute = 5/60 hour = 1/12 hour \[4 = \frac{d}{t+\frac{1}{12}}\]\[4t+\frac{1}{3} = d\] ''5 km/hr then he would have reached 10 minutes early.'' 10minutes = 10/60 hour = 1/6 hour \[5 = \frac{d}{t-\frac{1}{6}}\]\[5t - \frac{5}{6}= d\] \[5t - \frac{5}{6}=4t+\frac{1}{3}\]\[t=\frac{7}{6}hours\] \[d=4(\frac{7}{6})+\frac{1}{3} = 5 km\] Is it correct? (Even though I think it is more likely to be wrong :| )

OpenStudy (theviper):

@Callisto yes u r correct:) Thanx:)

OpenStudy (theviper):

@mathslover also